The alternator on a truck engine produces 350Wof electric power when it rotates at 4000 rev/min. If it is 95 percent efficient, what should the difference (in newtons) between the tensions in the tight and slack parts of its V-belt drive be? The diameter of the alternator’s pulley is 10 cm.  


1
Expert's answer
2021-11-05T18:19:19-0400
"\tau=\frac P\omega,\\\space\\
T=\frac \tau r=\frac{P}{\omega r},\\\space\\
\Delta T=T(1-\eta)=\frac{P}{\omega r}(1-\eta)=5.25\text{ N}."